OpenStudy (anonymous):

simplify: x^-1 + y^-1 / x^-2

OpenStudy (anonymous):

on the botton its x to the negative 2

OpenStudy (mertsj):

\[\frac{\frac{1}{x}+\frac{1}{y}}{\frac{1}{x^2}}\times\frac{x^2y}{x^2y}=\frac{xy+x^2}{y}\]

OpenStudy (anonymous):

i understand how you got that but the answer my teacher said was x squared

OpenStudy (mertsj):

Your teacher said the final answer is x^2?

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

So either your teacher is wrong or the problem is posted incorrectly. If you want to prove that it could not be x^2 choose numbers for x and y

OpenStudy (mertsj):

\[\frac{1}{x}+\frac{\frac{1}{y}}{\frac{1}{x^2}}\]

OpenStudy (anonymous):

no the fiirst prob i gave you exactly what i have on my paper

OpenStudy (mertsj):